in FRIEDEL-CRAFTS ACYLATION AND COLUMN CHROMATOGRAPHY lab

REACTION PROCEDURE: To a 5 mL conical vial, add a spin vane and connect a micro air condenser. Do not use one of the plastic caps to connect your conical vial to your air condenser. Make sure the drying tube is filled with drierite and plugged at both ends with cotton. The apparatus is sketched on next page drying tube - - micro air condenser clamp spin vane conical vial Add 0.150 g of aluminum chloride and 2.0 mL of methylene chloride to the conical vial. To prevent a significant amount of methylene chloride evaporating, replace the air condenser promptly after the addition is complete. Use a similar procedure to add the remaining reactants. Add an amount of acetyl chloride assigned by your TA of acetyl chloride (if using a glass pipette, assume that each drop is 0.01 mL). Stir the reaction mixture in an ice bath on a hot plate (no heating!). While the vial cools, prepare a solution of 0.100 g offerrocene in 0.5 to 1.0 mL of methylene chloride in a snap cap vial. Add this solution to the cooled vial. The mixture should quickly change from orange to a deep violet color. After replacing your condenser, remove the ice bath and allow the mixture to stir at room temperature While your reaction runs, add 4-6 mL of ice cold water to a 15 mL glass centrifuge tube. Once 15 minutes have elapsed, use a pipette to transfer your reaction mixture to the centrifuge tube containing ice water. Neutralize this mixture by adding 6M sodium hydroxide solution (about 0.5 mL). Use pH paper to confirm neutrality. If you add more NaOH than intended and the pH becomes basic, aluminum salts will precipitate, but this is OK, since the aluminum salts are essentially waste at this point. Make sure to mix the aqueous and non-aqueous phases at this point - it is essential to neutralize the acid in BOTH phases A pipette is the best method for mixing the phases. Once the reaction mixture has been neutralized, you will perform a microscale extraction (see PLKE Technique 12.4, p.722-726) to extract your ferrocene derivatives from the aqueous solution. After you separate the methylene chloride layer from the aqueous phase with a pipette, extract the aqueous phase twice more with 3 mL of methylene chloride each time. An emulsion often forms during the extraction, but luckily, centrifugation is one of the best ways to break an emulsion! If you do not have two discrete layers, briefly spin your sample in the centrifuge. Dont forget to balance the centrifuge with an extra tube (filled with the same solvents as your tube) or with the sample of another group labelled appropriately Combine the methylene chloride extracts in a 25 mL Erlenmeyer flask. Once you have combined your methylene chloride extracts, examine them. Are there drops of water present in the Erlenmeyer flask?They may be clinging to the walls of the flask or floating. If water droplets are visible, use a pipette to transfer your methylene chloride solution to a 50 mL Erlenmeyer flask

(2 pts) Amount of acetyl chloride used ____4______(drops) (each drop is 0.01 mL) _____2______(equivalents) Based on the amount of ferrocene you began the Friedel-Crafts reaction with, determine the percent yield of acetylferrocene, diacetylferrocene, and the percent recovery of unreacted ferrocene (show work)


react by 0.100 g of ferrocene at beginning

got 0 g of ferrocene, 0.012g of diacetylferrocene , 0.017g of acetylferrocene
Answer #1

calculate no of moles of each of them i.e ferrocene, acetylferrocene,diacetylferrocene. then divide the each of the mole value with moles of ferrocene taken and multiply by 100 we get %

for calculating no of moles u need molweight

mol weight of ferrocene=186.03

mol weight of acetyl ferrocene=228.07

mol weight of diacetyl ferrocene=270.10

amount of ferrocene taken for reaction=0.1g=0.1/186.03=5.37*10-4moles

amount of aceyl ferrocene formed= 0.017/228.07=7.45*10-5 moles

amount of diacetyl ferrocene formed= 4.44*10-5moles

% yield of diacetyl ferrocene= 8.27%

% yield of acetyl ferrocene= 13.87%